OpenStudy (anonymous):

**Calc 2 Help Needed** Find the length of the arc formed by y = (1/8) * (0.5x^2 -16ln(x)) from x=2 to x=6.

OpenStudy (zarkon):

what seems to be the problem?

OpenStudy (zarkon):

I assume you know the formula \[\int\limits_{a}^{b}\sqrt{1+(f'(x))^2}dx\]

OpenStudy (anonymous):

oh yes.. I do.. I can't figure out the derivative.

OpenStudy (anonymous):

of y = (1/8) * (0.5x^2 -16ln(x))

OpenStudy (zarkon):

\[\frac{dy}{dx}=\frac{1}{8}\left(x-\frac{16}{x}\right)\]

OpenStudy (anonymous):

Ok so now it is sqrt(1+ 1/8 * (x-16/x)

OpenStudy (zarkon):

no...look at the formula again.

OpenStudy (anonymous):

Ahh forgot ^2

OpenStudy (zarkon):

correct

OpenStudy (anonymous):

so 1+ ((x-16/8)^2

OpenStudy (zarkon):

no

OpenStudy (anonymous):

I can't bring the 1/8 in? Or shouldn't rather.

OpenStudy (zarkon):

you lost an x

OpenStudy (anonymous):

ohhh the ^2 wouldn't allow it.

OpenStudy (zarkon):

\[\frac{1}{8}\left(x-\frac{16}{x}\right)\] \[=\frac{1}{8}\left(\frac{x^2-16}{x}\right)\] \[=\left(\frac{x^2-16}{8x}\right)\] then 1+square it \[=1+\left(\frac{x^2-16}{8x}\right)^2\]

OpenStudy (anonymous):

so how did the top x-16 go to x^2-16?

OpenStudy (anonymous):

oh common denominator

OpenStudy (zarkon):

yes...\[x=\frac{x^2}{x}\] for x not equal to zero

OpenStudy (anonymous):

ok so after squaring I have 1+ ((x^2 - 16)^2/64x^2.. Well wolfram did.

OpenStudy (zarkon):

make into one fraction

OpenStudy (anonymous):

with the one?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

ok so now I have 64x^2 - (x^2 - 16)^2 all over 64x^2

OpenStudy (zarkon):

why the minus sign

OpenStudy (anonymous):

+ I mean.. sorry

OpenStudy (zarkon):

multiply out the numerator...then factor

OpenStudy (anonymous):

so now:( x^4 + 32x^2 + 256) / 64x^2

OpenStudy (zarkon):

yes...now factor the numerator

OpenStudy (anonymous):

the top would factor to (x^2 + 16)^2

OpenStudy (zarkon):

yes...now take the square root of the whole thing

OpenStudy (anonymous):

the sqrt would cancel the ^2

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

so (x+16)/64

OpenStudy (zarkon):

you lost an x again.

OpenStudy (zarkon):

also it is not 64 anymore

OpenStudy (anonymous):

Whoops.. (x+16)/8x?

OpenStudy (zarkon):

close \[\frac{x^2+16}{8x}\]

OpenStudy (anonymous):

Ohhh that one.. I am the king of mistakes tonight.

OpenStudy (zarkon):

write as \[\frac{x^2+16}{8x}=\frac{x^2}{8x}+\frac{16}{8x}=\frac{x}{8}+\frac{2}{x}\] and integrate from 2 to 6

OpenStudy (anonymous):

2.5

OpenStudy (zarkon):

no

OpenStudy (anonymous):

OMG why did I do that.. ok let me try again.

OpenStudy (anonymous):

I keep getting a negative.

OpenStudy (zarkon):

what did you get for the integral before evaluating the limits?

OpenStudy (anonymous):

Oh wow, I need to integrate..

OpenStudy (zarkon):

:)

OpenStudy (anonymous):

Sorry, you probably think I am absolutely dumb.

OpenStudy (zarkon):

no

OpenStudy (anonymous):

1.33333

OpenStudy (zarkon):

no

OpenStudy (anonymous):

After integrating I had x^2/16 + 4/2x

OpenStudy (zarkon):

\[\int\left(\frac{x}{8}+\frac{2}{x}\right)dx\] \[=\frac{x^2}{16}+2\ln|x|+c\]

OpenStudy (anonymous):

ok after wolfram it says the second part should contain a log

OpenStudy (anonymous):

4.19721 with a little confidence.

OpenStudy (zarkon):

yes...which equals \[2(\ln(3)+1)\]

OpenStudy (anonymous):

YESSS!!!!!! I can't believe that simple problem turned so nasty!

OpenStudy (zarkon):

;)

OpenStudy (anonymous):

Would you care to help me solve another?

OpenStudy (anonymous):

I'll start a new problem of course.

OpenStudy (zarkon):

hmmm...don't know if I have that much time. post it in a new thread...someone will help...If I have time I'll look

OpenStudy (anonymous):

Okay thanks so much.. Were you by any chance a math major?

OpenStudy (zarkon):

I'm a Math professor

OpenStudy (anonymous):

Cool, where at?

OpenStudy (zarkon):

so yes...I was a math major

OpenStudy (zarkon):

I don't give out that info :)

OpenStudy (anonymous):

Ahhh.. haha, I understand. Well I'm a student at a SEC school. Hopefully you aren't my teacher.

OpenStudy (zarkon):

I'm not in the SEC

OpenStudy (anonymous):

Haha okay.. Well thanks again for the help. This stuff is hard to understand in a lecture hall.

OpenStudy (zarkon):

good luck

OpenStudy (anonymous):

Thank you. Have a great night.